# Combining Multi-Objective Bayesian Optimization with Reinforcement Learning for TinyML

### v2

@@ 5. Evaluation
We applied our multi-objective DNN hyperparameter optimization on two use-cases with 150 samples each: (1) image classification (32x32 RGB images from CIFAR10 (Krizhevsky, 2012)) using a reduced version of ResNet18 (He et al., 2016) (three residual blocks instead of four, 1.6M initial parameters) and (2) time-series classification of (daily) human activities (DaLiAc) (Leutheuser et al., 2013) with a window length of 1024 datapoints, using a down-scaled version of MobileNetv3 (Howard et al., 2019) (2.8M initial parameters). Compared to models commonly used by relevant benchmarks for embedded AI, e.g., TinyML Perf (Banbury et al., 2021), we did not reduce the initial numbers parameters as our optimization process uses pruning and therefore dynamically adjusts the exact number of neurons during training. We trained our models using Stochastic Gradient Decent (SGD) with an exponential learning rate decay. For the optimization, we chose 256 KB of RAM, 1 MB of ROM and $1e^{9}$ FLOPS as constraints imposed by common microcontroller platforms.
The detailed list of hyperparameters considered by our optimizer and their sampling intervals derived from expert knowledge can be found in Appendix A. We consider seven parameters controlling general properties of DNN training (training length, batch size, properties of the optimizer and learning rate) and four that control the filter pruning. We use the automated gradual pruning algorithm (Zhu and Gupta, 2017) to generate our pruning schedule. It takes a start and end epoch calculated relatively to the number of training epochs and a number of iterative steps as an input. Our optimization considers the pruning sparsity hyperparameter, i.e., how many filters are removed, for each convolutional layer separately. Therefore, the search space considered increases with the depth of the optimized DNN as for each prunable layer an additional pruning sparsity parameter is added which means that the search space is extended by one dimension.
Figure 1. ARS-MOBOpt compared to baseline approaches (top row: MobileNetv3 on DaLiAc; bottom row: ResNet on CIFAR10). Left column: Our approach (ARS-MOBOpt) outperformed all other optimization. Remaining columns: feasible Pareto sets proposed by our RL-based MOBOpt of DNNs (blue) compared to other Bayesian and evolutionary approaches (in both cases we find samples dominating those proposed by the other algorithms while meeting imposed target constraints).
@@ 5.1. Algorithmic Benchmark
As our first experiment we compared the quality of our approach (ARS-MOBOpt) to several well-known Bayesian optimization strategies (ParEGO (Knowles, 2006), TurBO (Eriksson et al., 2019), MorBO (Daulton et al., 2022)), one evolutionary approach (NSGA-II (Deb et al., 2002)), and a random sampling approach as a baseline, see Fig. 1. All Bayesian optimization strategies were implemented using BoTorch (Balandat et al., 2020). We gave each optimization strategy a search budget of 150 samples to evaluate the objective functions and monitored the improvement of the feasible hypervolume over the samples. For all tested Bayesian strategies the first 10 samples were conducted as priors using Latin Hypercube sampling (McKay et al., 2000). We normalized our objectives accuracy, ROM and RAM consumption, and FLOPs to their respective constraints and used the maximum feasible objective value as a common reference when computing the hypervolume. We performed 5 independent seeds for each strategy and provide the observed variance as well.
Fig. 1 shows the results for two common DNN architectures, MobileNetv3 and ResNet18, which we trained using two datasets, DaLiAc and CIFAR10. The plots show the observed hypervolume over the number of optimization samples. Therefore, the resulting curves, one for each optimization strategy, describe the improvement of the hypervolume. First, all informed solvers outperformed the independent random sampling baseline on both use-cases. Second, all the approaches that are based on Bayesian optimization, i.e., TurBO, MorBO, ParEGO and ARS-MPBOpt, which consider the constraints of a very tight search budget of 150 samples, performed better then the evolutionary sampling strategy (NSGA-II). Additionally, for both use-cases a significant Hypervolume improvement can be observed when using our ARS-based sampling strategy with 3000 sampling directions, a top-$k$ selection percentage of 1%, and a rollout horizon of 4 to train the competing policies. This implies that by using a trainable RL-based strategy to determine new promising parametrizations as the optimization progresses, our algorithm makes much better use of the knowledge about the search space encoded in the surrogate model than the other non-trainable Bayesian sampling approaches.
@@
To get more insight into the general performance of ARS, we tested it on a subset of the multivariate datasets of the UEA & UCR time series classification benchmark (Bagnall et al., 2017) and compared ourselves to ParEGO using the same setup as described above, see Table 1 for all data sets on which reasonable models with a predictive accuracy of at least 70% are found (and Table 4 in Appendix B for a complete listing). To provide context, we also list reference results for InceptionTime (Ismail Fawaz et al., 2020) that we extracted from (Bagnall et al., 2017). We let both ARS and ParEGO optimize a CNN architecture with a single regular convolution followed by 10 depth-wise separable convolutions with batchnorm and with a total of 5 squeeze and excitation blocks (Hu et al., 2018) after convolutions 10, 8, 6, 4, and 2, followed by adaptive average pooling and a single linear layer for classification (2.6m initial parameters). Our approach achieves a better hypervolume than ParEGO in 7 of 10 data sets (and is en par on one data set). If we also consider data sets where we found models with predictive accuracies below 70% our approach outperforms ParEGO in approx. two-thirds of the data sets, see Appendix B. From the results we also see that compared to ParEGO, our algorithm was generally better at finding DNNs with higher accuracy scores, but often at the cost of higher memory and FLOPs requirements.
@@ 5.2. Robustness and Hyperparameter Selection
Our proposed ARS sampling strategy can be configured with a number of hyperparameters that affect the behavior of the ARS algorithm, see Alg. 1. The quality of the results obtained by our algorithm can vary depending on the parameterization, see Fig. 2: In general, choosing a learning rate of $1e^{-3}$, an exploration noise of $1e^{-2}$, a top-$k$ selection rate of $1\%$, a number of sampling directions around $3000$, and a rollout horizon of $4$ yielded the best results. In the following, we want to further discuss the impact of key hyperparameters, i.e., the number of sampling directions $N$, the top-$k$ percentage of directions selected for policy updates, and the number of steps $H$ taken for each episode (rollout horizon), to provide a broader understanding of their effects.
@@
First, especially for complex optimization problems with a high-dimensional parameter space, a high number of sampling directions significantly improved the hypervolume that our ARS-based sampling strategy was able to achieve, compared to a low number of directions, which often resulted in a below-average performance, see Fig. 2(a). However, we also found, that it is not feasible to increase the number of directions indefinitely to support more and more complex problems, as this also significantly increases the time required to train the ARS policies and the memory consumption of our algorithm. We found that the exact number of sampling directions which is required to obtain a competitive result is highly correlated with the complexity of the underlying optimization problem.
@@
Second, when increasing the top-$k$ percentage of sampling directions selected by ARS to perform policy updates, we observed that only selecting a small subset of $1\%$ of the sampled directions to perform policy updates improved the hypervolume achieved by ARS compared to using larger subsets, see Fig. 2(b). Our observation is thus consistent with (Mania et al., 2018)’s argument for introducing the hyperparameter as part of ARS. The authors point out that including all observed rewards in the parameter update can be detrimental to performance because outliers can easily introduce bias.
@@
Third, we found that an optimal episode length (rollout horizon) is at $H=4$ steps, see Fig. 2(c). Both a significantly smaller and larger episode length on average resulted in a decreased performance. Still, we want to point out that we observed a significantly larger variance for longer episode lengths compared to shorter episode lengths. This is consistent with the increased variance in the estimation of total expected return in MC-like algorithms like ARS (Sutton and Barto, 2018).
5.3. Comparison with MCUNet and Zero-Shot NAS

### v3

@@ 5. Evaluation
Parameter Search Space Interval
epochs uniform $[100,500]$
batch size uniform $[20,200]$
learning rate (lr) log. uniform $[1e^{-5},1e^{-2}]$
momentum log. uniform $[0.7,0.99]$
lr schedule uniform $[0.4,0.9]$
lr gamma uniform $[0.4,0.9]$
weight decay log. uniform $[0.6,0.99]$
pruning start uniform $[0.0,0.6]$
pruning end uniform $[0.8,0.95]$
pruning steps uniform $[1,20]$
pruning sparsity uniform $[0.1,0.99]$
We applied our multi-objective DNN hyperparameter optimization to two use cases with 150 samples each: (1) image classification (32x32 RGB images from CIFAR10 (Krizhevsky, 2012)) using a reduced version of ResNet18 (He et al., 2016) (three residual blocks instead of four, 1.6M initial parameters) and (2) time-series classification of (daily) human activities (DaLiAc) (Leutheuser et al., 2013) with a window length of 1024 datapoints, using a down-scaled version of MobileNetv3 (Howard et al., 2019) (2.8M initial parameters). Compared to models commonly used by relevant benchmarks for embedded AI, e.g., TinyML Perf (Banbury et al., 2021), we did not reduce the initial numbers parameters as our optimization process uses pruning and therefore dynamically adjusts the exact number of neurons during training. We trained our models using Stochastic Gradient Decent (SGD) with an exponential learning rate decay. For the optimization, we chose 256 KB of RAM, 1 MB of ROM and $1e^{9}$ FLOPS as constraints imposed by common microcontroller platforms.
The detailed list of hyperparameters considered by our optimizer and their search space intervals derived from expert knowledge can be found in Table 1. These parameters also define the action space of our ARS agents. We consider seven parameters that control general properties of DNN training (training length, batch size, optimizer properties, and learning rate) and four that control filter pruning. We use the automated gradual pruning algorithm (Zhu and Gupta, 2017) to generate our pruning schedule. t takes as input a start and end epoch calculated relative to the number of training epochs and a number of iteration steps. Our optimization considers the pruning sparsity hyperparameter, i.e., how many filters are removed, for each convolutional layer separately. Therefore, the search space considered increases with the depth of the optimized DNN as for each prunable layer an additional pruning sparsity parameter is added which means that the search space is extended by one dimension.
@@ 5.1. Algorithmic Benchmark
As a first experiment, we compared the quality of our approach (ARS-MOBOpt) with several well-known Bayesian optimization strategies (ParEGO (Knowles, 2006), TurBO (Eriksson et al., 2019), MorBO (Daulton et al., 2022)), one evolutionary approach (NSGA-II (Deb et al., 2002)), and a random sampling approach as a baseline, see Fig. 2. All Bayesian optimization strategies including our novel ARS-based BO strategy were implemented with BoTorch (Balandat et al., 2020), which we used in conjunction with Optuna (Akiba et al., 2019) as a metaheuristic optimization framework. For NSGA-II and Random we used the implementations provided by Optuna.
We gave each optimization strategy a search budget of 150 samples to evaluate the objective functions and monitored the improvement of the Hypervolume over the samples. Higher Hypervolume values are considered superior. For all tested Bayesian strategies the first 10 samples were conducted as priors using Latin Hypercube sampling (McKay et al., 2000). We normalized our objectives accuracy, ROM, RAM, and FLOPs to their respective constraints and used the maximum feasible objective value as a common reference when computing the Hypervolume. We performed five independent seeds for each strategy and provide the observed variance.
Figure 2. ARS-MOBOpt compared to several baseline approaches (top row: MobileNetv3 on DaLiAc; bottom row: ResNet on CIFAR10). Left column: Our approach (ARS-MOBOpt) outperforming all others in terms of Hypervolume after 50 to 75 evaluated samples. Remaining columns: feasible Pareto sets determined by ARS-MOBOpt compared to other Bayesian (ParEGO, TurBO, MorBO) and evolutionary (NSGA-II) approaches with random sampling (Random) as a baseline.
[…]
ARS-MOBOpt compared to five baseline approaches TurBO, MorBO, qParEGO, NSGA-II, and random sampling. The figure has two rows and two columns of plots. The top row shows results for MobileNetv3 on DaLiAc, and the bottom row shows results for ResNet on CIFAR10. The plots in the left column show that our approach (ARS-MOBOpt) outperforms all other baselines in terms of Hypervolume after 50 to 75 evaluated samples. The plots in the right column show the feasible Pareto sets determined by ARS-MOBOpt compared to the baselines.
Fig. 2 shows the results for two common DNN architectures, MobileNetv3 and ResNet18, which we trained using two datasets, DaLiAc and CIFAR10. The plots show the observed Hypervolume over the number of optimization samples. Therefore, the resulting curves, one for each optimization strategy, describe the improvement of the Hypervolume. First, all informed solvers outperformed the independent random sampling baseline on both use cases. Second, all the approaches that are based on Bayesian optimization, i.e., TurBO, MorBO, ParEGO and ARS-MPBOpt, which consider the constraints of a very tight search budget of 150 samples, performed better than the evolutionary sampling strategy (NSGA-II). Additionally, for both use cases, a significant Hypervolume improvement is observed when using our ARS-based sampling strategy with 3000 sampling directions, a top-$k$ selection percentage of 1%, and a rollout horizon of 4 to train the competing policies. This implies that by using a trainable RL-based strategy to determine new promising parametrizations as the optimization progresses, our algorithm makes much better use of the knowledge about the search space encoded in the surrogate model than the other non-trainable Bayesian sampling approaches.
@@
To gain more insight into the overall performance of ARS, we tested it on a subset of the multivariate datasets from the UEA & UCR time series classification benchmark (Bagnall et al., 2017). We compare our method with ParEGO using the same setup as described above, see Table 2 for all datasets on which reasonable models with an accuracy of at least 70% are found (and Table 4 in Appendix A for a complete list). To provide context, we also list reference results for InceptionTime (Ismail Fawaz et al., 2020) that we extracted from Bagnall et al. (2017). We let both ARS and ParEGO optimize a CNN architecture with a single regular convolution followed by 10 depth-wise separable convolutions with batchnorm and with a total of 5 squeeze and excitation blocks (Hu et al., 2018) after convolutions 10, 8, 6, 4, and 2, followed by adaptive average pooling and a single linear layer for classification (2.6m initial parameters). Our approach achieves a better Hypervolume than ParEGO in 7 of 10 datasets (and is en par on one dataset). If we also consider datasets where we found models with predictive accuracies below 70% our approach outperforms ParEGO in almost two-thirds of the datasets, see Appendix A. The results also show that, compared to ParEGO, our algorithm was generally better at finding DNNs with higher accuracy scores, but often at the cost of higher memory and FLOPs requirements.
@@ 5.2. Robustness and Hyperparameter Selection
[…]
Detailed results examining the effects of the three most important ARS hyperparameters, sampling directions, top-k selected samples, and rollout horizon, on the achieved Hypervolume for MobileNetv3 trained on the DaLiAc dataset with a window length of 1024. The three plots in the figure show that 3000 sample directions, a top-k selection rate of 1%, and a rollout horizon of 16 resulted in the best Hypervolume improvement over 150 samples.
@@
Our proposed ARS sampling strategy can be configured with different hyperparameters that affect the behavior of the ARS algorithm, see Alg. 1 for a complete listing. The most important are (a) the sampling directions $N$, (b) the top-$k$ selection rate, and (c) the rollout horizon $H$. Both the optimization progress and the quality of explored solutions obtained by our algorithm can vary depending on their parameterization, see Fig. 3(c): In general, a top-$k$ selection rate of $1\%$, a number of sampling directions around $3000$, and a rollout horizon of $4$ yielded the best results. In the following, we want to further discuss the impact of these key hyperparameters, to provide a better understanding of their effects.
@@
First, especially for complex optimization problems with a high-dimensional parameter space, a high number of sampling directions significantly improved the Hypervolume that our ARS-based sampling strategy was able to achieve, compared to a low number of directions, which often resulted in a below-average performance, see Fig. 3(a). However, we found, that it is not feasible to increase the number of directions indefinitely to support more and more complex problems, as this also significantly increases the time required to train the ARS policies and the memory requirements of our algorithm. Instead, it should be noted that the exact number of sampling directions required to obtain a competitive result is highly correlated with the complexity of the underlying optimization problem.
@@
Second, when increasing the top-$k$ percentage of sampling directions selected by ARS to perform policy updates, we observed that only selecting a small subset of $1\%$ of the sampled directions to perform policy updates improved the Hypervolume achieved by ARS compared to using larger subsets, see Fig. 3(b). Our observation is thus consistent with Mania et al. (2018)’s argument for introducing the hyperparameter as part of ARS. The authors point out that including all observed rewards in the parameter update can be detrimental to performance because outliers can easily introduce bias.
@@
Third, we found that an optimal episode length (rollout horizon) is at $H=4$ steps, see Fig. 3(c). Both significantly shorter and longer episodes resulted in lower average performance. However, we would like to point out that we observed a significantly larger variance for longer episodes compared to shorter episodes. This is consistent with the increased variance in the estimation of total expected return in MC-like algorithms such as ARS (Sutton and Barto, 1999).
5.3. Discussion
To better understand why our RL-based solver can improve on other Bayesian solvers like ParEgo, we analyzed both using a synthetic optimization problem. Given a two-dimensional single-objective optimization problem with an objective landscape as shown in Fig. 7 from  Passino (2005) in Appendix B. This problem has a challenging landscape with large areas that have little to no gradient combined with some steep minima and maxima. The respective Bayesian surrogates of ParEgo and ARS-MOBOpt as well as the topographies of their expected improvement (EI) after 40 trials are shown in Figs. 4(a) and 4(b). For our approach (ARS-MOBOpt), we also show the rollouts performed during the validation of the local policies for five selected start points (red crosses) after the last (40th) sample.
We manually selected two sets of priors (marked as red triangles) to provide both algorithms with interesting starting situations: First, a situation where all samples are close together in a region with almost no gradient, see top left in Fig. 4(a) and 4(b). This initialization is interesting because it provides almost no information about the environment, requiring extensive exploration by the algorithms before they can exploit found minima. Second, a more general starting situation with a broader distribution of prior samples, see top right in Fig. 4(a) and 4(b). This initialization focuses more on evaluating how well the algorithms can apply the exploration-exploitation tradeoff and how fast they can find minima. For both priors, we took care not to include samples close to the global minimum.
(a) ParEGO
(b) ARS
Figure 4. Topography of the optimization landscape (top row) and EI (bottom row) estimated by the Bayesian surrogates for the synthetic optimization problem shown in Fig. 4(b) in Appendix B after 40 samples for both ParEGO and ARS (ours), given two sets of priors marked with red triangles. The global minimum is at $\theta_{0}=15$, $\theta_{1}=5$. For ARS, the rollouts of the trained policies of the 40th sample are shown as lines with their starting points marked by red crosses.
[…]
The topography of the optimization landscape and the EI estimated by the Bayesian surrogates for the synthetic optimization problem after 40 samples for both ParEGO and ARS, given two sets of priors marked by red triangles. For ARS, the rollouts of the trained policies of the 40th sample are shown as lines with their starting points marked by red crosses. The results in the figures show that after 40 samples, ARS was able to sample closer to the global optimum of the synthetic problem than ParEGO. Furthermore, the rollouts of the trained policies show that the ARS agents have learned to navigate the already explored optimization landscape in search of a minimum.
Looking at the results proposed by ParEgo (marked as black dots) after 40 samples in Fig. 4(a), it becomes evident that for both sets of priors the algorithm seems to have a clear focus on exploration, leading to a wide sampling of the search space. However, looking at the two examples, this can lead to situations where even though the global optimum was either not found (left column in Fig. 4(a)) or it was not fully exploited (right column in Fig. 4(a)). Furthermore, we argue that this focus on exploration is mainly a result of the definition of the EI as it does not consider any ”negative improvement”. Looking at the topography of the EI for both examples in the last row of Fig. 4(a), it becomes evident that this can result in large parts of the acquisition function having no non-zero gradient. We argue that this is not optimal when relying on gradient-based solvers, as the ParEgo implementation typically does.
In comparison, if we look at the result of our proposed ARS-based approach in Fig. 4(b), it can be seen that our competing local policies, visualized as lines with red crosses as their starting point, have become experts in solving the environment around their starting point and can therefore provide sound solutions even in regions without significant gradients. Therefore, compared to ParEgo, we observed a more directed exploration and a clearer focus on exploiting environmental knowledge, compared to Fig. 4(a). However, despite this more pronounced focus on exploitation, due to the competing ensemble of agents, our solver was still able to consistently escape local minima, e.g. see Fig. 4(b) (right col.).
5.4. Comparison with Related Optimization Approaches
In the following, we provide comparative results of our approach with other RL, NAS, and automated pruning methods. Namely, we analyze the feasibility of using a different type of RL agent, PPO, instead of ARS, and compare our approach to both MCUNet, zero-shot NAS, and automated pruning.
5.4.1. Comparison with PPO
The following experiment is used to compare the performance of ARS with PPO (Schulman et al., 2017). PPO is an on-policy gradient RL algorithm that, unlike comparable algorithms such as TRPO (Schulman et al., 2015), requires only first-order optimization while achieving better overall stability and performance. Because PPO is sample-efficient and can be used for environments with continuous action spaces, we selected it as a suitable candidate for comparison with the ARS agents we have used so far.
We repeated the optimization of ResNet on CIFAR10 and MobileNetv3 on DaLiAc as described in Section 5.1 with PPO instead of ARS and provide results in Fig. 5(c). For context and comparison, we also show the results for ARS and random sampling which we took from Fig. 2. For all experiments, we used the stablebaselines (Raffin et al., 2021) implementation of PPO and repeated each 5 times, recording means and standard deviations as we did with ARS and random sampling in Section 5.1.
(a) PPO Hyperparameter pre-optimization for ResNet on CIFAR10
(b) ResNet on CIFAR10
(c) MobileNetv3 on DaLiAc
Figure 5. Comparison of MOBOpt with ARS and PPO for ResNet on CIFAR10 and MobileNetv3 on DaLiAc. For both cases we show random sampling as the baseline. The results for ARS and Random are the same as the results in Fig. 2. Since the default hyperparameters of PPO did not produce results of quality better than random sampling for ResNet, we performed a hyperparameter pre-optimization over 20 trials, see (a). The Hypervolume results are shown in (b) for ARS-MOBOpt and pre-optimized PPO. (c) shows the need of problem-specific pre-optimization of hyperparameters for PPO. The green curve is obtained when applying the pre-optimized parameter settings from (a) to MobileNetv3 rather than ResNet.
[…]
Comparison of MOBOpt with ARS and PPO for ResNet on CIFAR10 and MobileNetv3 on DaLiAc. For both cases, random sampling is shown as the baseline. In (a), it can be seen that different PPO hyperparameters resulted in a high variance of the Hypervolume in the pre-optimization for ResNet on CIFAR10. In (b), we show that using the default hyperparameters of PPO results in a Hypervolume that is no better than random sampling, while PPO with optimized hyperparameters from the pre-optimization in (a) results in a Hypervolume that is as good as when using ARS. (c) shows that using PPO with the same hyperparameters optimized for ResNet on CIFAR10 in (a) does not work for MobileNetv3 on DaLiAc, i.e. the achieved Hypervolume is not better than random sampling, demonstrating the need for problem-specific pre-optimization of hyperparameters for PPO.
Initially, we tested PPO with its default parameters. However, this resulted in a worse Hypervolume improvement than random sampling, see the orange and red curves in Fig. 5(b). Furthermore, when analyzing the orange curve, it can be observed that during the first part of the optimization PPO performed well in efficiently exploring the search space, resulting in a roughly similar Hypervolume compared to ARS after about 40 samples. However, the improvement started to stagnate sharply after this initial exploration-heavy phase. As a result, we found that after the initial exploration phase, all PPO agents started to have a strong tendency to get trapped in local optima found by previous samples, which resulted in the agents completely stopping their exploration.
Considering the above results, we performed a PPO hyperparameter optimization using Optuna (Akiba et al., 2019) for the ResNet on CIFAR10 optimization problem, see Fig. 5(a). For each PPO configuration tested, we ran MOBOpt on 20 samples with a fixed prior of 2 samples. A complete listing of the search space can be found in Appendix C. The results presented in Fig. 5(a) show that simply changing the hyperparameters of PPO, while leaving the MOBOpt problem as it is, led to very different Hypervolume improvements, i.e., drastically different behaviours of the PPO agents. Most notably, we observed that for many configurations, PPO was not able to improve on the prior at all, while for other configurations, the algorithm achieved a high Hypervolume and a significant improvement over the prior.
Using the best performing set of PPO hyperparameters from the optimization shown in Fig. 5(a), we repeated the experiment for ResNet on CIFAR10, see the green curve in Fig. 5(b). It can be observed that by using the optimized PPO configuration the algorithm achieved a Hypervolume improvement that consistently even exceeded the one we observed for ARS. Moreover, especially during the first 40 samples, PPO was able to improve the Hypervolume much faster than any of the other algorithms we tested.
Finally, we tested how well the PPO hyperparameters optimized for ResNet on CIFAR10 abstracts to another related optimization problem. Therefore, we repeated the second experiment we discussed earlier in Fig. 2, i.e. MobileNetv3 on DaLiAc, with PPO and the hyperparameters optimized for ResNet, see Fig. 5(c). The results of this additional experiment show that while PPO was again initially able to reach a better Hypervolume faster than all other algorithms we tested, after the first 20 samples the improvement sharply stagnated and the algorithm stopped finding significantly better points. This is similar to the observations we made when testing PPO for ResNet on CIFAR10 with default parameters. As a result, with the hyperparameters from Fig. 5(a) optimized for ResNet, PPO again barely outperformed the random sampling baseline for the MobileNetv3 on DaLiAc optimization problem, and even performed worse than when using its default parameters, compare green and orange curves in Fig. 5(c).
Overall, the results we obtained align with common observations for deep reinforcement learning algorithms like PPO, where optimal hyperparameters can vary significantly across tasks, often necessitating extensive tuning for new tasks. We have shown that the performance of PPO and its exploration-exploitation tradeoff can be significantly improved by performing a full study of its hyperparameters to adapt them as best as possible to a specific optimization problem. However, especially compared to ARS, where we have discussed the effects of its three hyperparameters on optimization in detail in Section 5.2, and where we have shown that they work consistently well for a wide range of tasks, PPO has significantly more complex hyperparameters whose effects are not intuitive, and which we have shown to behave differently for different MOBOpt problems. We conclude that an effective use of PPO would necessitate a full optimization of its hyperparameters for each optimization problem, making it an extremely time-consuming and resource-intensive optimizer for MOBOpt. We emphasize this point because one of the main reasons we chose ARS in the first place was its simplicity.
5.4.2. Comparison with MCUNet and Zero-Shot NAS